At 27°C density of a gaseous substance at 3 bar is same as that of hydrogen at 9 bar. What is the molar mass of the substance?

Grade:11

1 Answers

Arun
6 years ago
d1=m1p1/Rt and d2=m2p2/Rt here temperature is constant then d1=m1p1 and d2=m2p2 (m1=? m2=1 P1=3 P2=9) M1=M2p2/p1= m1=1×9/3=m1=3g/mol